OpenStudy (anonymous):

Solve the problem by writing an inequality. A club decides to sell T-shirts for $15 as a fundraiser. It costs $20 plus $9 per T-shirt to make the T-shirts. Write and solve an equation to find how many T-shirts the club needs to make and sell in order to profit at least $150. 15x – (9x + 20) ≥ 150; x ≥ 28.33 15x – 9x + 20 ≥ 150; x ≥ 20 (8x + 20) – 15x ≥ 150; x ≥ 20 15x – 9(x + 20) ≥ 150; x ≥ 20

OpenStudy (anonymous):

ok,let x be the number of t-shirts to sell so total selling price (SP) = 15x now lets calculate cost price (CP) it costs 20+9 per tshirt--->20+9x profit = SP-CP and atleast means >= 150 so final inequality is 15x-(20+9x)>=150

OpenStudy (anonymous):

A